# Non-equilibrium Quantum Monte Carlo Algorithm for Stabilizer Rényi Entropy in Spin Systems

### v3

@@ Appendix C Implementation of SSE on Ising Model - Zero Temperature
To illustrate this point, let us consider a simple example. At $h=0$, after initial thermalization, the initial product states are $|00\cdots 0\rangle$ or $|11\cdots 1\rangle$ with equal probabilities, and the operators that flip the spins ($H_{k,0}$) will not show up in the operator lists. Consequently, SSE only gives the configurations corresponding to the mixed states $\rho=(|00\cdots 0\rangle\langle 00\cdots 0|+|11\cdots 1\rangle\langle 11\cdots 1|)/2$, which gives $M_{2}(\rho)=\ln 2$, while either $|00\cdots 0\rangle$ or $|11\cdots 1\rangle$ gives $M_{2}(\rho)=0$ as pure states. Because of the ground degeneracy, for a pure state $|g\rangle=\alpha|00\cdots 0\rangle+\beta|11\cdots 1\rangle$, SRE varies with choice of coefficients $\alpha$ and $\beta$. To resolve these issues, we fix the initial product state as $|00\cdots 0\rangle$, giving an unambiguous evaluation of SRE. Note that the above discussions also applies to the case of $h\ll 1$.
For ground state SRE, one needs to use sufficiently large length of operator list $m$ to ensure the convergence towards the exact ground state SRE. In Fig. 7(a), we show the SRE density at the chosen $m\in\{9600/10000,12800,16000\}$, where we observe that the SRE are well converged, despite that for some of them (e.g., $N=400$ at $h=3.10$ and $h=3.30$), the SRE decrease slightly from $m=10000$ to $m=16000$.
In Fig. 7(b), we show the SRE density versus $h$, and observe that SRE densities intersect at different $h$ at larger system size from that at smaller system size. In Fig. 7(c), we directly compare our results with those from Ref. [27] which are in good agreement except for at $h=2.9$ at large $L$.
@@
Figure 7: (a) Convergence of 2nd SRE density on ground state of 2D transverse field Ising model (PBC, $J=1$, $N=400$ with dot markers, $N=64$ with cross markers), w.r.t. the length of operator list $m$. (b) 2nd SRE density on ground state of 2D transverse field Ising model (PBC, $J=1$) versus the strength of field $h$. (c) Comparison between the QMC data (dot markers) and the data from Ref. [27] (square markers).
Appendix D Connection of Lattice Spins at the End of Non-equilibrium Process
@@
Appendix E Matrix Product State Based Methods
@@
Appendix G Statistical Properties of the Non-equilibrium QMC Algorithm
@@
Figure 10: Numerical examination on Eqn. 32, y-axis is the variance $\sigma^{2}$ directly evaluated from the sample set $W$, the fitting parameters $\alpha_{c}$ and $\alpha$ comes from the fitting on the SNR of $e^{-W}$ (a) $T=0$, $\alpha\simeq 1.20$, $\alpha_{c}\simeq 2.48\cdot 10^{-3}$; (b) $T>0$, $\alpha\simeq 1.20$, $\alpha_{c}\simeq 6.11\cdot 10^{-4}$. $y=x$ (black dotted lines) are drawn in (a) and (b) for comparisons.
@@
Appendix H A Unifying Framework for Exponential Observable Calculation
@@
Appendix I SRE of Mixed States

### v4

@@ III.2 Implementation of SSE on Ising Model - Zero Temperature
To illustrate this point, let us consider a simple example. At $h=0$, after initial thermalization, the initial product states are $|00\cdots 0\rangle$ or $|11\cdots 1\rangle$ with equal probabilities, and the operators that flip the spins ($H_{k,0}$) will not show up in the operator lists. Consequently, SSE only gives the configurations corresponding to the mixed states $\rho=(|00\cdots 0\rangle\langle 00\cdots 0|+|11\cdots 1\rangle\langle 11\cdots 1|)/2$, which gives $\widetilde{M}_{2}(\rho)=\ln 2$, while either $|00\cdots 0\rangle$ or $|11\cdots 1\rangle$ gives $\widetilde{M}_{2}(\rho)=0$ as pure states. Because of the ground degeneracy, for a pure state $|g\rangle=\alpha|00\cdots 0\rangle+\beta|11\cdots 1\rangle$, SRE varies with the choice of coefficients $\alpha$ and $\beta$. To resolve these issues, we fix the initial product state as $|00\cdots 0\rangle$, giving an unambiguous evaluation of SRE. Note that the above discussions also applies to the case of $h\ll 1$.
@@ III.3 Numerical Results
Figure 4: (a) Convergence of 2nd SRE density on ground state of 2D transverse field Ising model (PBC, $J=1$, $N=400$ with dot markers, $N=64$ with cross markers), w.r.t. the length of operator list $m$. (b) 2nd SRE density on ground state of 2D transverse field Ising model (PBC, $J=1$) versus the strength of field $h$. (c) Comparison between the QMC data (dot markers) and the data from Ref. [27] using TTN (square markers).
For ground state SRE, one needs to use sufficiently large length of operator list $m$ to ensure the convergence towards the exact ground state SRE. In Fig. 4(a), we show the SRE density at the chosen $m\in\{9600,10000,12800,16000\}$, where we observe that the SRE are well converged, despite that for some of them (e.g., $N=400$ at $h=3.10$ and $h=3.30$), the SRE decrease slightly from $m=10000$ to $m=16000$. In Fig. 4(b), we show the SRE density $\widetilde{m}_{2}\equiv\widetilde{M}_{2}/N$ versus $h$, and observe that SRE densities intersect at different $h$ at larger system size from that at smaller system size. In Fig. 4(c), we directly compare our results with those from Ref. [27] using tree tensor networks (TTN), which are in good agreement except for at $h=2.9$ at large $L$.
For the rest of the numerical experiments, we choose $m$ between $10000$ and $16000$, and use 160 path samples (i.e. the number of non-equilibrium work iterations) for ground state calculations and 640 for finite temperature calculations for all the system sizes considered here. We choose $\Delta\lambda=10^{-4}$. For a single path-sample on our largest system of $N=16\times 16$ computing the ground state SRE takes $\lesssim 4$ CPU hours using a code implemented in C++.
In Fig. 5, we show the comparisons between the 2nd SRE density $\widetilde{m}_{2}$ from the non-equilibrium QMC and those from matrix product states (MPS) in various scenarios: (a) ground state SRE of 1D Ising ring (i.e., periodic boundary conditions, PBC); (b) ground state SRE of 2D Ising cylinder (i.e., PBC on one spatial direction and open boundary condition on the other, PBC-OBC); (c) finite-$T$ SRE of 2D Ising cylinder. They all display quantitative agreements when comparing against with the results from the tensor-network methods. Remarkably, in case (c), the noticeable discrepancies between the QMC and MPS data result from the errors of limited bond dimension $\chi$ and the finite Trotter splitting time interval $\tau$ for the MPS method as can be seen from the observation that the MPS data converges towards the QMC data as we increase $\chi$ and decrease $\tau$. This suggests that to reach the same accuracy as QMC, the MPS-based methods requires more computational resources and careful extrapolation. It is interesting to observe from Fig. 5 that SRE density peaks at the phase transition for the ground states of the 1D transverse-field Ising model, but not in 2D cases. It is an interesting open question to understand this observation.
In Fig. 6, we report our results on larger systems, and we are able to draw the following conclusions: (1) 2nd SRE mainly grows linearly with system size for short-range correlated pure states, as indicated by the ground states on 2D transverse-field Ising model. (2) At $T>0$, there exist subleading terms from 2nd SRE weakly dependent on $T$, as evident from the crossing in Fig. 6(b). In Fig. 7, we show the 2nd SRE versus a broader range of temperature, and find two extremals at $\beta\simeq 0.36$ and $0.78$ for sufficiently large systems. It is unclear whether the $\beta$ of these extremals have any particular physical significance.
IV Computational Cost
For $T=0$ or ground state SRE, we use projector QMC and the key difference for the time complexity is that the length of the operator list $m$ has to be sufficiently large to decrease the contributions from the excited states and guarantee the convergence to the ground state. We argue that the lower bound for $m$ such that the relative bias error to the exact ground state SRE $\widetilde{M}_{n}(g)=M_{n}(g)$ is below $\delta_{r}$ is given by (see SM for the derivation)
V Conclusion
In this work, we develop a novel algorithm to compute the SRE of finite temperature and ground state systems in all dimensions for sign-problem free Hamiltonians. To accomplish this, we show how to contract a tensor network representation of the SRE using non-equilibrium QMC. It has proven to be accurate and efficient, regardless of the spatial dimensions and temperature of the model, as observed from the numerical experiments on the transverse-field Ising model and the complexity analysis of this algorithm. We expect this algorithm would serve as a useful numerical tool to understand quantum systems from the new perspective of quantum magic. Interesting future directions include, but are not limited to, an in-depth study of phase transition features from $n(\geq 2)$th SRE consisting of the derivatives [48], the application to models with deconfined quantum critical points [60, 61, 62, 63, 64] or field theory [21, 65, 66], the extension of this approach to Hamiltonians with sign problems, and the exploration of replacing the non-equilibrium QMC algorithm with other efficient algorithms for evaluating the exponential observable [37, 38, 35].
Acknowledgements.
We would like to thank T.C. Lu, Z.Y. Meng, J. Zhao, J. D’Emidio, M. Dalmonte, E. Tirrito, M. Frau and P. Tarabunga for insightful discussions and valuable comments. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and which is supported by funds from the University of Illinois at Urbana-Champaign. We acknowledge support from the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Award No. 2016136).
Appendix A Connection of Lattice Spins at the End of Non-equilibrium Process
@@
Appendix B Matrix Product State Based Methods
@@
Appendix D Statistical Properties of the Non-equilibrium QMC Algorithm
@@
Figure 10: Numerical examination on Eqn. 34, y-axis is the variance $\sigma^{2}$ directly evaluated from the sample set $W$, the fitting parameters $\alpha_{c}$ and $\alpha$ comes from the fitting on the SNR of $e^{-W}$ (a) $T=0$, $\alpha\simeq 1.20$, $\alpha_{c}\simeq 2.48\cdot 10^{-3}$; (b) $T>0$, $\alpha\simeq 1.20$, $\alpha_{c}\simeq 6.11\cdot 10^{-4}$. $y=x$ (black dotted lines) are drawn in (a) and (b) for comparisons.
@@
Appendix E A Unifying Framework for Exponential Observable Calculation
@@
Appendix F SRE of Mixed States